# Consistency in Language Models: Current Landscape, Challenges, and Future Directions

### v1

@@ 3.2 Multilingual Consistency
Similarly to other aspects of NLP research, the overwhelming majority of studies on foundation model consistency are currently performed on English-language data, significantly limiting our understanding of model consistency. To broaden this understanding, more research is needed on both monolingual consistency in non-English languages and on cross-language consistency behaviors.
3.2.1 Consistency in Non-english Languages
There exists a sizable gap between the available amount of training data for English and that available for all other languages Üstün et al. (2024). While more than 7,000 languages are spoken around the world today, an astounding 73% of popular datasets used to train LLMs are primarily or entirely English Longpre et al. (2023). This severe sampling bias in the construction of datasets results in disparities in model performance across languages even on well-studied tasks Lai et al. (2023).
@@ 3.2.1 Consistency in Non-english Languages
Moreover, inherent differences between languages may significantly influence the consistency of language models trained on them. Structural features such as word order or inflectional morphology can vary in their stability across languages Dediu & Cysouw (2013). These differences can make it more difficult to train models that produce consistent output for certain languages, even when all languages are equally represented in the training data. More research is necessary to understand the effect of linguistic differences and limitations of multilingual training data on consistency in non-English languages.
3.2.2 Cross-lingual Consistency
@@
Cross-lingual consistency refers to whether a model produces compatible or equivalent output when the same query is presented in different languages. Recent work has demonstrated significant challenges in this area. Shen et al. (2024) found that LLMs exhibit inconsistent safety behaviors across languages, with safety guardrails being more easily circumvented in non-English languages. Similarly, Xing et al. (2024) observed that LLMs produce inconsistent factual information when queried about the same knowledge in different languages, suggesting knowledge representation gaps across languages. Qi et al. (2023) examined factual consistency across languages and identified disparities, with languages more dissimilar to English less likely to reflect synthetically inserted factual associations through model editing. Jin et al. (2023) evaluated cross-lingual inconsistency specifically in healthcare questions, finding discrepancies in medical advice across languages. Zhou & Zhang (2024) explored how political biases manifest inconsistently in bilingual models, revealing that models may express different political positions depending on the input language.
@@
3.3 Automatic and Human Evaluation
Most previous studies have used automatic metrics alone to assess the level of consistency in language models. Although automatic evaluation is necessary to ensure objectivity and fast assessment of large data sizes, human evaluation is important to establish an acceptable baseline. It becomes especially important when model consistency is assessed in the context of highly sensitive or subjective culture-based applications (e.g. social appropriateness) or when the automatically measured consistency level is sufficiently high.
@@ 3.3 Automatic and Human Evaluation
The evaluation of consistency poses several unique challenges. Automatic metrics often struggle to capture the full spectrum of consistency types (factual, logical, semantic), while human evaluation suffers from subjectivity and cognitive biases. The contextual nature of consistency requires evaluation across multiple responses, different phrasings, and various contexts, making comprehensive assessment computationally expensive and logistically challenging. Further complicating matters, consistency evaluation interacts with other dimensions such as factuality, helpfulness, and safety—a model may be internally consistent but factually incorrect, or it may sacrifice consistency to maintain safety.
Interestingly, there is no publicly available benchmark yet to measure all different aspects of consistency in language models across diverse tasks. Future work should focus on developing high-quality benchmarks that address the above challenges.
3.4 The Harms and Benefits of Consistency
@@
Inconsistent outputs from language models can lead users to receiving conflicting or incorrect information, which is particularly problematic in scenarios where factual accuracy is crucial (Tam et al., 2023; Wang et al., 2024)—such as in medical, legal, or financial contexts—especially when such information is used for decision-making. In critical systems, such as autonomous vehicles or medical diagnosis support, inconsistent responses can lead to critical safety risks. In less critical applications, inconsistent responses lead to poor user experience, causing frustration and reducing overall utility of the system (Lazar et al., 2023; van Bergen et al., 2024; Zhang et al., 2024a).
In addition to the above problems, inconsistent responses can reflect and magnify the underlying societal biases and stereotypes in training data, leading to potentially discriminatory outcomes for certain user groups, amplifying unfair use of the models, and causing representational harm, including demeaning social groups (Blodgett et al., 2020).
@@
While inconsistency causes certain risks and potential harms to the users of language models in real-world applications, it is also important to understand its potential value and advantages in specific settings and situations. First, some degree of inconsistency can lead to more diverse and creative output, which can be valuable in tasks requiring originality or brainstorming. As such, inconsistency might reflect the ability of a model to adapt to different contexts or user needs, potentially providing more personalized responses.
Second, inconsistent outputs of a model prompt users to engage more critically with the content and seek additional verification. This can be beneficial in educational applications, as long as the level of possible model inconsistency is known and carefully controlled.
Finally, inconsistencies can help identify areas where the model needs improvement, driving further research and development in AI.
3.5 Methods to Improve Consistency
@@
Despite the importance, surprisingly few approaches have emerged that actually increase the consistency of language models. Current proposals to do so fall into two narrow categories. The first uses fine-tuning techniques to improve consistency between multiple generations from a language model when supplied with same or similar outputs. To this end, (Elazar et al., 2021) used a custom loss function, Raj et al. (2025) used knowledge distillation from more consistent teacher models, and Raj et al. (2025); Zhao et al. (2024) utilized synthetic datasets of groups of consistent input-outputs. The second approach attempts to improve self-consistency, i.e. consistency between the reasoning process a model goes through and the final answer it produces Deng et al. (2023); Wang et al. (2023); Wei et al. (2022).
@@
While these methods show promise, they primarily address symptoms rather than the fundamental causes of inconsistency. There remains a critical need for research investigating the structural basis of consistency in language models’ representational spaces, consistency-oriented pretraining objectives, and architectures designed to maintain consistency across diverse contexts. Such foundational approaches could potentially eliminate the trade-offs between consistency and other valuable properties like creativity or adaptability that the above methods tend to exhibit.
4 Conclusion and Call to Action
In this paper, we provide a review of consistency in language models, highlighting challenges, current approaches, and future directions in this critical area of research. We have shown that despite significant advances, the field lacks standardized definitions, evaluation metrics, benchmarks, and mitigation techniques. This lack of standardization hampers progress and makes it difficult to compare the results between studies.
@@ 4 Conclusion and Call to Action
We specifically call on the research community to address several key challenges: (1) developing standardized definitions and taxonomies of consistency types; (2) creating comprehensive, multilingual, and cross-lingual benchmarks for consistency evaluation; (3) establishing robust evaluation protocols that combine automatic metrics with human evaluation; (4) investigating the relationship between consistency and other important properties such as factuality, safety, and helpfulness; and (5) developing efficient methods to enhance consistency without sacrificing other beneficial model capabilities.
To this end, we emphasize the need for interdisciplinary collaboration, bringing together perspectives from linguistics, psychology, philosophy, and ethics to better understand the multifaceted nature of consistency in human and machine language use. By addressing these challenges collectively, we can move toward language models that exhibit more reliable, trustworthy, and human-aligned behavior across diverse contexts and applications.
@@ Appendix A Appendix: Multimodal Consistency
Until 2022, every consistency study was analyzing robustness of language models to various text perturbations or to semantically equivalent texts only. Starting in 2022, some interest in non-textual modalities started to appear that comes primarily from text-to-image model analysis. For example, Tan et al. (2022) explores the challenge of generating consistent and high-quality images from given texts in the task of visual-language understanding and highlights the need to design a better text-image consistency metric, a problem that remains under-explored in the community. In their study, Tan et al. (2022) present a novel CLIP-based metric named Semantic Similarity Distance (SSD) that leads to significantly better text-image consistency while maintaining decent image quality. The attempts to quantify the consistency in text-to-image models are continued by Berglund et al. (2024), which proposes a novel semantic consistency score for image generation that has strong agreement with human annotators. Recently, there was an attempt to evaluate the understanding capability of generative models in both language and vision domains West et al. (2024). West et al. (2024) conducted interrogative evaluation of image understanding models via visual question answering in an open-ended setting. They investigated whether the models produce consistent output when interrogated about the content of their generated image, and figured out that although models can outperform humans in generation, they regularly show evidence of inconsistency between their generation and understanding performance.

### v2

@@ 3 Discussion and Recommendations
Similarly to other topics in NLP research, the overwhelming majority of studies on LM consistency are English-based, significantly limiting our understanding of the topic. To broaden this understanding, more research is needed on both monolingual consistency in non-English languages and on cross-language consistency behaviors.
@@
There is a substantial gap between the amount of training data available for English and that available for all other languages Üstün et al. (2024). While more than 7,000 languages are spoken around the world today, an astounding 73% of the popular datasets used to train LLMs are primarily or entirely English Longpre et al. (2023). This severe sampling bias in dataset construction results in disparities in model performance between languages, even in well-studied tasks Lai et al. (2023). Inherent differences between languages may also significantly influence the consistency of the LMs trained on them. Structural features such as word order or inflectional morphology can vary in their stability across languages Dediu & Cysouw (2013). These differences can make it more difficult to train models to produce consistent output for certain languages, even when all languages are equally represented in the training data. More research is necessary to understand the effect of linguistic differences and limitations of multilingual training data on consistency in non-English languages.
@@
Recent work has demonstrated significant challenges in cross-lingual consistency, i.e., whether a model produces compatible or equivalent outputs when the same query is presented in different languages. Shen et al. (2024) found that LLMs exhibit inconsistent safety behaviors across languages, with safety guardrails being more easily circumvented in non-English languages. Xing et al. (2024) observed that LLMs produce inconsistent factual information when asked about the same knowledge in different languages, suggesting knowledge representation gaps across languages. Qi et al. (2023) examined factual consistency across languages and found that languages more dissimilar to English are less likely to reflect synthetically inserted factual associations through model editing. Jin et al. (2023) evaluated cross-lingual inconsistency specifically in healthcare questions and found discrepancies in medical advice across languages. Zhou & Zhang (2024) explored how political biases manifest inconsistently in bilingual models, revealing that models may express different political positions depending on the input language.
@@
Consistency Evaluation
@@
Evaluating consistency has several unique challenges. Most previous studies have used automatic metrics alone to assess consistency in LMs. Although automatic evaluation can ensure objectivity and fast assessment, human evaluation is important to establish an acceptable baseline, especially in highly sensitive or subjective culture-specific applications (e.g. social appropriateness), or when automatic metrics are sufficiently high. Automatic metrics often struggle to capture the different nuances of consistency (factual, logical, semantic), while human evaluation suffers from subjectivity and cognitive biases. The contextual nature of consistency requires evaluation across multiple responses, different phrasings, and various contexts, making comprehensive assessment computationally expensive and logistically challenging. Further complicating matters, consistency evaluation interacts with other dimensions such as factuality, helpfulness, and safety—a model may be internally consistent but factually incorrect, or it may sacrifice consistency to maintain safety.
While several consistency benchmarks have recently emerged, there remains a need for more comprehensive evaluation frameworks that measure all different aspects of consistency in LMs across diverse tasks. Recent benchmarks have made important contributions but typically focus on specific consistency types or limited task domains (Jang et al., 2022; Bonagiri et al., 2024; Cui et al., 2024; Liu et al., 2024b; Paleka et al., 2024; Liu et al., 2024a; Gilhuly & Shahzad, 2025). Future work should focus on developing more holistic benchmarks that address the breadth of consistency challenges outlined above.
Impact
@@
Inconsistent output can cause users of language-based systems to receive conflicting or incorrect information. This is problematic in scenarios where factual accuracy is crucial (Tam et al., 2023; Wang et al., 2024)—such as in medical, legal, or financial contexts—especially when such information is used for decision making. In critical systems, such as autonomous vehicles or medical diagnosis support, inconsistent responses can lead to critical safety risks. In less critical applications, inconsistent responses lead to poor user experience, cause frustration, and reduce overall utility (Lazar et al., 2023; van Bergen et al., 2024; Zhang et al., 2024a). Inconsistency can also reflect and magnify the underlying societal biases and stereotypes in the training data, leading to potentially discriminatory outcomes for certain user groups, amplifying unfair use and causing representational harm (Blodgett et al., 2020).
@@
Inconsistency may have some advantages in specific situations. Lower degrees of consistency can lead to diverse and creative outputs, which can be valuable in tasks requiring originality or brainstorming. Inconsistency might reflect the ability of a model to adapt to different contexts or user needs, potentially providing more personalized responses. Inconsistent outputs of a model prompt users to engage more critically with generated content, avoid overreliance, and seek additional verification. This can be beneficial in educational applications, provided the level of possible inconsistency is carefully calibrated.
Improving Consistency
@@
There are surprisingly few approaches that actually increase the consistency of LMs. Current proposals to do so fall into two narrow categories. The first approach employs fine-tuning to improve consistency between multiple generations from a LM when supplied with the same or similar inputs. Elazar et al. (2021) used a custom loss function, Raj et al. (2025) used knowledge distillation from more consistent teacher models, and Raj et al. (2025); Zhao et al. (2024) used synthetic datasets of groups of consistent input-outputs. The second approach attempts to improve self-consistency, i.e. consistency between a model’s reasoning process and the final answer Deng et al. (2023); Wang et al. (2023); Wei et al. (2022).
@@
Albeit promising, the above methods primarily address symptoms rather than the fundamental causes of inconsistency. There remains a critical need for research investigating the structural basis of consistency in LMs’ representational spaces, consistency-oriented pre-training, and architectures designed to maintain consistency across diverse contexts. Such foundational approaches may eliminate the trade-offs between consistency and other valuable properties like creativity and diversity.
4 Call to Action
@@
We call on the research community to address several key challenges: (1) developing standardized definitions and taxonomies of consistency types; (2) creating comprehensive, multilingual, and cross-lingual benchmarks for consistency evaluation; (3) establishing robust evaluation protocols that combine automatic metrics with human evaluation; (4) investigating the relationship between consistency and other important properties such as factuality, safety, and helpfulness; and (5) developing efficient methods to enhance consistency without sacrificing other beneficial model capabilities. To this end, we emphasize the need for interdisciplinary collaboration, bringing together perspectives from linguistics, psychology, philosophy, and ethics to better understand the multifaceted nature of consistency in human and machine language use. By addressing these challenges collectively, we can move toward LMs that exhibit more reliable, trustworthy, and human-aligned behavior across diverse contexts and applications.
@@ Appendix A Appendix: Multimodal Consistency
Until 2022, every consistency study was analyzing robustness of LMs to various text perturbations or to semantically equivalent texts only. Starting in 2022, some interest in non-textual modalities started to appear that comes primarily from text-to-image model analysis. For example, Tan et al. (2022) explores the challenge of generating consistent and high-quality images from given texts in the task of visual-language understanding and highlights the need to design a better text-image consistency metric, a problem that remains under-explored in the community. In their study, Tan et al. (2022) present a novel CLIP-based metric named Semantic Similarity Distance (SSD) that leads to significantly better text-image consistency while maintaining decent image quality. The attempts to quantify the consistency in text-to-image models are continued by Berglund et al. (2024), which proposes a novel semantic consistency score for image generation that has strong agreement with human annotators. Recently, there was an attempt to evaluate the understanding capability of generative models in both language and vision domains West et al. (2024). West et al. (2024) conducted interrogative evaluation of image understanding models via visual question answering in an open-ended setting. They investigated whether the models produce consistent output when interrogated about the content of their generated image, and figured out that although models can outperform humans in generation, they regularly show evidence of inconsistency between their generation and understanding performance.